OpenStudy (anonymous):

GIVING MEDALS f and f^-1 are functions. if f(2)=pi+7, then f^-1(pi+7)=?

OpenStudy (anonymous):

@Preetha @nincompoop

OpenStudy (gorv):

\[f*f^{-1}=1\]

OpenStudy (gorv):

multip-ly both side by f^{-1}

OpenStudy (gorv):

can u do it @em2000

OpenStudy (anonymous):

i dont know how

OpenStudy (gorv):

\[f^{-1} *f(2)=f^{-1}(\pi+7)\]

OpenStudy (gorv):

\[f^{-1}*f=1\]

OpenStudy (gorv):

on left side we will left with only 2

OpenStudy (gorv):

so \[f^{-1} (\pi+7)=???\]

OpenStudy (gorv):

@em2000

OpenStudy (anonymous):

10.14??

OpenStudy (gorv):

on left side f inverse and f will cancel each other so we left with 2 only

OpenStudy (anonymous):

OMG I GET IT

OpenStudy (anonymous):

THANK U SO MUCH SORRY IT JUST TOO A WHLE @gorv

OpenStudy (gorv):

you r welcome :)